OpenStudy (anonymous):

find the average value of the function over the given interval. f(x)=1/x , [1,e]

OpenStudy (nikvist):

\[<f(x)>=\frac{1}{e-1}\int\limits_{1}^{e}\frac{dx}{x}=\frac{1}{e-1}\ln{x}|_{1}^e=\frac{1}{e-1}(1-0)=\frac{1}{e-1}\]